OpenStudy (anonymous):

Find the value of 7 + 32 (-5 + 1) ÷ 2.

OpenStudy (anonymous):

do the brackets first so 32x-4

OpenStudy (anonymous):

I got 17.5

OpenStudy (anonymous):

32 multiplied by -4

OpenStudy (anonymous):

7 + 32 = 39. (-5 + 1) = -4 39 x -4 = 35 35 / 2 = 17.5

OpenStudy (anonymous):

Using PEMDAS, we first get rid of the parenthesis:\[7+32(-4)\div 2\]next we multiply:\[7-128\div2\]then divide:\[7-64\]and lastly we subtract:\[-57\]